OpenStudy (anonymous):

What is the average rate of change of f(x), represented by the graph, over the interval [-1, 1]?

OpenStudy (anonymous):

ill see if I can get the graph

OpenStudy (anonymous):



OpenStudy (anonymous):

that's what it looked like

OpenStudy (anonymous):

Is that y = 1/x

OpenStudy (anonymous):

its just a standard graph the answers to it are 2 3 5 and 6 the graphs numbered to 11

OpenStudy (anonymous):

I utterly have no idea what im doing I just came back to school from injury so im totally lost on the topic

OpenStudy (anonymous):

We need to know the y-values at -1 and 1 Average rate of change is basically the slope between the two points that's why we need the y-coordinates

OpenStudy (anonymous):

I guess if you can just estimate them from the graph that will give us a close enough answer

OpenStudy (anonymous):

i know what slope is i just don't know how to put it in this type of thing what would be your best guess i think its 5

OpenStudy (anonymous):

I can't really tell without knowing the coordinates. Like we have (-1, \(y_1\)) and (1, \(y_2\)). We need to know \(y_1\) and \(y_2\) to get slope. \[\frac{ f(b)-f(a) }{ b-a }\] is just another way of writing \[\frac{ y_2-y_1 }{ x_2-x_1 }\] It's the same thing, we just need to know the y's to solve. Can you post a screenshot or photo of the graph?

OpenStudy (anonymous):